# supplementary

### v3

@@ 3.1. Connection to Classical Hypothesis Testing
We first briefly address testing for contrasts in Bayesian linear regression without spatial effects, i.e., $\gamma=0$ in (1). Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ non-zero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003). We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are the generalized least squares estimates. We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate posterior probabilities for the absolute standardized linear combinations among elements of $\beta$ to exceed a threshold $\epsilon$. Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
@@
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 00. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
@@
for any given $\epsilon$. Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We draw exact posterior samples from the Normal-Gamma family. Specific domain applications may provide further insights into possible constraints on $\epsilon$ and what is considered to be a clinically meaningful difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean outcomes for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
Classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 00, while small $v_{k}(\epsilon)$ suggests $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 00. In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities yield identical decisions on contrasts in the following sense.
@@
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$, and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value for testing the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
@@
See LABEL:append:prop1_proof.
Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
@@ 3.2. Extension to Spatial Random Effects
We consider (1), where each observation corresponds to one region. We let $V_{\phi}$ be known, whereas $\rho$ can either be a fixed hyperparameter or a random variable with a prior density. Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we conduct posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$. For $k=1\ldots,K$, we define
@@
In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the rankings of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ are stable with respect to $\epsilon$ for any $\rho\in(0,1)$.

### v4

@@ 3.1. Connection to Classical Hypothesis Testing
We first briefly address contrast testing in Bayesian linear regression without spatial effects, that is, $\gamma=0$ in (1). Whereas $\mathbb{P}(c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0\,|\,y)=0$ for a continuous posterior distribution, we can test $H^{(k)}_{0}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H^{(k)}_{1}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $K$ nonzero $p\times 1$ vectors $c_{1},\ldots,c_{K}$ using a classical linear model, $y=X\beta+\eta,\eta\sim N_{n}(0_{n},\sigma^{2}V_{y})$, where $\beta$ is fixed but unknown (e.g., seberHypothesisTesting2003). We compute $t_{\text{obs}}^{(k)}=\frac{c_{k}^{\mathrm{\scriptscriptstyle T}}\widehat{\beta}}{\widehat{\sigma}\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}c_{k}}}$ for each $c_{k}$, where $\widehat{\beta}=(X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$ and $(n-p)\widehat{\sigma}^{2}=\left(y-X\widehat{\beta}\right)^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}\left(y-X\widehat{\beta}\right)$ are generalized least squares estimates. We reject $H^{(k)}_{0}$ if the p-values $p_{k}=\mathbb{P}\left(T>\left|t^{(k)}_{\text{obs}}\right|\right)$, where $T\sim t_{n-p}$, are below a multiplicity adjusted threshold.
For Bayesian inference, we retain the continuous conjugate prior on $\beta$ and evaluate the posterior probabilities for the absolute standardized linear combinations between elements of $\beta$ to exceed a threshold $\epsilon$. Hence, for all $k=1,\ldots,K$ and any $\epsilon>0$, we define
@@
We refer to these posterior probabilities as $\epsilon$-difference probabilities, where $\epsilon$ represents a universal threshold across all $K$ linear combinations indicating the number of posterior standard deviation units that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is away from 0. If $\beta^{(1)},\ldots,\beta^{(N)}$ and $\sigma^{(1)},\ldots,\sigma^{(N)}$ are $N$ samples drawn from their respective posterior distributions, then we estimate (6) using
@@
for any given $\epsilon$. Adapting (2) to $\gamma=0$ yields $\pi(\beta,\sigma^{2}\,|\,y)=N(\beta\,|\,Mm,\sigma^{2}M)\times\text{IG}(\sigma^{2}\,|\,a,b)$, where $a=a_{0}+\frac{n}{2}$, $b=b_{0}+\frac{1}{2}\left(y^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m^{\mathrm{\scriptscriptstyle T}}Mm\right)$, $M^{-1}=M_{0}^{-1}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}X$ and $m=m_{0}+X^{\mathrm{\scriptscriptstyle T}}V_{y}^{-1}y$. We draw exact posterior samples from the Normal-Gamma family. Specific domain applications may provide further information on possible constraints on $\epsilon$ and what is considered to be a clinically significant difference between coefficients $\beta_{i}$ and $\beta_{j}$, which, for example, could represent mean results for different treatment groups in a Bayesian Analysis Of Variance (ANOVA).
The classical p-values and the probabilities $\left\{v_{k}(\epsilon)\right\}_{k}$ defined in (6) have opposite interpretation: a small value of $p_{k}$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is far from 0, while small $v_{k}(\epsilon)$ suggests that $|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|$ is likely to be close to 0. In fact, Bayesian inference with a flat prior on $\beta$ amounts to using the classical generalized least squares estimate $\widehat{\beta}$. Hence, with a specific prior, the p-values and $\epsilon$-difference probabilities yield identical decisions on contrasts in the following sense.
@@
Let $c_{1},\ldots,c_{K}\in\mathbb{R}^{p}$, $X\in\mathbb{R}^{n\times p}$, $y\in\mathbb{R}^{n}$ and $V_{y}\in\mathbb{R}^{n\times n}$ be fixed. Under the standard linear model $y\sim N_{n}(X\beta,\sigma^{2}V_{y})$, let $p_{k}$ be the p-value to test the null hypothesis $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ against $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ for $k=1,\ldots,K$. Then, $p_{k}>p_{k^{\prime}}$ if and only if $v_{k}(\epsilon)<v_{k^{\prime}}(\epsilon)$ for all $k\neq k^{\prime}$, where $v_{k}(\epsilon)=\mathbb{P}\left(\left.\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\beta|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}Mc_{k}}}>\epsilon\;\right|\;y\right)$ under $\pi(\beta,\sigma^{2})\propto\pi(\sigma^{2})$.
@@
See Appendix LABEL:append:prop1_proof.
Theorem 1 implies that the ascending order of the p-values associated with null hypotheses of the form $H_{0}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta=0$ with the two-sided alternative $H_{1}^{(k)}:c_{k}^{\mathrm{\scriptscriptstyle T}}\beta\neq 0$ is equivalent to the descending order of the Bayesian $\epsilon$-difference probabilities. Since this result holds for all $\epsilon>0$ and the classical procedure does not depend on $\epsilon$, the Bayesian rankings are stable with respect to $\epsilon$.
@@ 3.2. Extension to Spatial Random Effects
We consider (1), where each observation corresponds to one region. We let $V_{\phi}$ be known, while $\rho$ can be either a fixed hyperparameter or a random variable with a prior density. Generalizing the $\epsilon$-difference probabilities in the fixed effects model, we perform posterior inference on random variables $r_{k}(\epsilon)=\text{I}\left(\frac{|c_{k}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}|}{\sigma\sqrt{c_{k}^{\mathrm{\scriptscriptstyle T}}M_{\star}c_{k}}}>\epsilon\right)$ for any given $\epsilon>0$. For $k=1\ldots,K$, we define
@@
In the context of detecting spatial disparities, consider $K$ pairs of neighboring regions $L=\{(i,j):i<j,i\sim j\}$, where $i\sim j$ means that the regions $i$ and $j$ are neighbors. If $c_{ij}\in\mathbb{R}^{n+p}$ and $a_{ij}\in\mathbb{R}^{n}$ are such that $c_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma_{\star}=a_{ij}^{\mathrm{\scriptscriptstyle T}}\gamma=\gamma_{i}-\gamma_{j}$, then the corresponding $\epsilon$-difference probability is $h_{ij}(\epsilon;\rho)=\mathbb{P}\left(\left.\frac{|\phi_{i}-\phi_{j}|}{\sqrt{a_{ij}^{\mathrm{\scriptscriptstyle T}}\operatorname{Var}(\phi\,|\,y,\sigma^{2},\rho)a_{ij}}}>\epsilon\;\right|\;y,\rho\right)$. Using (1) and (2), we extend the insights obtained from Theorem 1 and show that the ranking of conditional posterior probabilities $h_{1}(\epsilon),\ldots,h_{K}(\epsilon)$ is stable with respect to $\epsilon$ for any $\rho\in(0,1)$.